

********************************************************************
UARS WEEKLY STATUS REPORT - 05 AUG 2005
********************************************************************
***********
SOLSTICE
***********
This week SOLSTICE was in day only mode until Monday evening at
05/213-23:43 when it was powered off completely. As usual, the
instrument worked wonderfully until it's shut down, taking solar data
on all available orbital tracking periods.
***********
SUSIM
***********
As far as is determined at this time, SUSIM has successfully
completed its complete set of lamp and reference scans from July 30
through August 1. Earlier, daily mid (1.1 nm) resolution, MgII, and
continuous monitoring scans for time periods with UARS telemetry or
tape recording were successfully performed.
SUSIM's final data was gathered on August 1. No more operations are
planned.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005079 has been processed and is waiting for archival
at the DAAC. HEPSA data for 2005080 to 2005089 is currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005089 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB processing code has been generated
and is currently under testing.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 July 2005 (Orbit 75889, GMT Day 204) through 29
July 2005 (Orbit 75993, GMT Day 210).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 207/1353 75942 FULLTIME TO DAY-ONLY
HALOE 208/0412 75951 POWER OFF
The beta angle ranged between the angles of 37.2 degrees to 34.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 4 degrees on 09 August 2005 (DOY 221).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 6.23 to 5.85 degrees C. Additional
battery performance data for each battery on 22 July 2005 and 29 July
2005 are provided in Appendix D.
The clock error ranged from 05.1 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
SOLSTICE = 01
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
One (1) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On Monday 11 July 2005 the results of the senior review at NASA HQ
were received. The direction given is that UARS will fly with HALOE
observations only until the end of the calendar year, followed by the
decommissioning of the UARS spacecraft in the new year. Final
calibrations for SUSIM were executed on 30 July and 01 August 2005.
SOLSTICE will perform its final calibration on 04 August 2005.
The UARS OBC clock rollover effort continued this week with the final
testing and simulation runs of the uplink procedure. All entities are
prepared for the uplink and ready to transition their systems to
support the new epoch year. The uplink is scheduled for 02 August on
the 214/1409-1435 event, with preliminary setup commanding on the
214/1307 event. All impacted facilities should take note of the
uplink times and be prepared for the epoch change.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 09 August 2005 (DOY 221).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 July 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75901 204/2059 05-222 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
75942 207/1353 05-223 SOLSTICE FULLTIME TO DAY-ONLY
75951 208/0412 05-225 MPS VT LEVEL 5
75951 208/0412 05-224 HALOE POWER OFF
75956 208/1212 05-226 MPS VT LEVEL 6
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 July 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 July 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75901 204/2101 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 July 2005
22 July 2005 - GMT Day 203 Beta = 35.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 67.2/ 39.2 -2.42/-3.68 +00.0/+00.0 05.8 0.00 0.0
2 67.2/-100.8 +10.06/+8.73 +31.6/-20.8 25.8 1.04 24.2
3 156.8/-700.0 +4.50/+3.55 +19.2/-11.6 25.8 1.05 13.7
29 July 2005 - GMT Day 209 Beta = 34.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 336.0/ 224.0 -2.42/-3.68 +00.0/+00.0 05.4 0.00 0.0
2 -117.6/-268.8 +10.73/+9.72 +28.8/-16.0 26.7 1.09 19.9
3 22.4/-677.6 +4.50/+3.87 +19.2/-10.4 26.7 1.06 13.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
75951 208/0412 LV6 -> LV5
75956 208/1212 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 July 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT 75916 205/2056 30832 Generic Late Acquisition
05-032 FOT 75942 207/1326 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-036 FOT 75831 200/0440 N/A ECLIPSE Command Hung-up
05-037 FOT 75859 202/0216 40172 Late Acq: bad weather @
Guam
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT See List July 05 30832 Generic Late Acquisition